Bronz Snails company hired you as a consultant to estimate the

company’s WACC .

The firm’s target capital structure is 30.5% Debt, 13.1% Preferred stock and 56.4%

Common Equity.

The Firms noncallable bonds mature in 15years. The bonds have a 9.5% annual

coupon rate, a par value of $1,000 and a market price of $1,135. Bonds pay coupon

payments semi annually. The firm has 200,000 bonds outstanding.

The firm has 7%, $100 par value preferred stocks. There are 1M shares outstanding.

The preferred stock currently sells at $98 per share.

Common Equity investors bond yield risk premium 6%.

Risk free rate is 2.25%, market risk premium 10.5% and common stock beta is 1.15.

Company’s tax rate 30%.

The firm has 20M shares outstanding of common stocks that sell at $21 per share.

The firm just paid a dividend of $1.80 per share and the constant growth rate is

expected to be 5.5%

The firm would like to use the lowest of the three methods (Bond yield risk

premium, CAPM and DCF) to estimate the cost of equity and it doesn’t expect to

issue new common stock.

Calculate the WACC? Do not round your intermediate calculations.
Answer #1

WACC = Cost of debt * weight of debt +Cost of equity * weight of equity+Cost of Pref.equity * weight of Pref.equity

Total value of Capital (V) = Market value of ( debt+pref equity+ Equity)

V=200000 *1135+ 1000000*98+ 20000000*21 = 745,000,000

weight of debt = 200000 *1135/745,000,000 = 0.305

weight of pref equity=1000000*98/745,000,000 =0.131

weight of equity=20000000*21/745,000,000= 0.564

Cost of debt = 9.5%*(1-30%)= 6.65%

Cost of pref equity= 7%

Cost of equity

  • Bond Yield method , Re = 9.5%+ 6% = 15.5%
  • CAPM , Re= 2.25%+1.15*10.5% = 14.33%
  • Dividend discount model , 21= 1.80 (1+0.055)/(Re- 0.055)

Re= 14.54%

Minimum equity cost in CAPM model i.e. 14.33%

WACC = 6.65%*0.305+7%*0.131+14.33%*0.564 = 11.03%